# Low-energy $N\phi$ scattering from a pole-enhanced triangle diagram

### v1

@@ III Results and discussion
Kinematical consistency requires the cutoff $\Lambda$ in the loop integral to be larger than the typical momentum flowing through the kaon lines. For a small cutoff value, $\Lambda=120\,\mathrm{MeV}$, noticeable deviations from the uncertainty bands are observed. In contrast, for $\Lambda=300$-$500\,\mathrm{MeV}$ the real part of the $N\phi$ scattering length remains stable and overlaps with the bands, supporting the relevance of the triangle mechanism in the physical-mass case.
@@
Figure 8: The real part of the $N\phi$ scattering length from Two-Kaon-Exchange with physical masses, where the solid, dashed, and dotted lines correspond to the central value, lower and upper limits of $\delta$, respectively. The vertical dash-dot-dot line represents the physical $\phi$ mass with respect to $K\bar{K}$ threshold.
@@
Such threshold-enhanced triangle dynamics is not unique to the $N\phi$ system and has been widely discussed in heavy exotic hadrons. Representative examples include the role of $D\bar{D}$ rescattering in $D\bar{D}\pi$ loops for the $X(3872)$ Dai et al. (2020); Yan et al. (2022); Cao et al. (2024), the contribution of $\Sigma_{c}\bar{D}\pi$ loops to positive-parity $P_{c}(4457)$ states Peng et al. (2021a); Wu et al. (2024b), and the importance of $DD\pi$ loops in the description of the $T_{cc}$ state Fleming et al. (2021); Yan and Valderrama (2022); Jia et al. (2023a, b); Dai et al. (2023). In these systems, the proximity of two- and three-body thresholds leads to non-perturbative enhancements analogous to those identified here.
A similar mechanism is expected to operate in channels where $D\bar{D}\pi$ couples to $Z_{c}(4020)$ or $X(4020)$, promoted by the nearby $X(3872)$ or $W_{c1}(3872)$ poles Peng et al. (2023a); Ji et al. (2025). In such cases, the triangle diagram generates an interaction that plays a role comparable to that of scalar-meson exchange, supplementing the axial-vector exchange mechanism Yan et al. (2021). Related threshold effects have also been discussed for hidden-charm and hidden-strangeness pentaquark states, where the $P_{c}(4457)$ and $P_{cs}(4560)$ states are enhanced by lower-lying $P_{c}(4312)$ and $P_{cs}(4338)$ poles through $\bar{D}\Sigma_{c}\pi$ and $\bar{D}\Xi_{c}\pi$ loops, respectively Peng et al. (2023b); Du et al. (2020); Fernández-Ramírez et al. (2019); Wang and Wang (2020); Liu et al. (2019); Wang et al. (2020); Chen et al. (2021); Xiao et al. (2021); Yan et al. (2023); Nakamura and Wu (2023); Zhu et al. (2023).
@@
In the dibaryon sector, a closely related dynamical mechanism has been invoked to interpret the $d(2380)$ as a $\Delta\Delta$ molecular state, where the existence of a near-threshold $N\Delta$ bound state plays a crucial role Gal and Garcilazo (2013). A similar picture also emerges in the $IJ=21$ and $12$ $N\Delta$ scattering channels when poles in the $NN$ subsystem are taken into account Gal and Garcilazo (2014). However, the situation for the $d(2380)$ is considerably more involved: the resonance is located between the $N\Delta\pi$ and $\Delta\Delta$ thresholds, and the large intrinsic width of the $\Delta$ renders a clean separation between two-body and three-body dynamics highly nontrivial Peng et al. (2022). This contrasts with the $N\phi$ system, where the thresholds are much better separated and the triangle mechanism can be isolated more transparently.
Another instructive comparison is provided by the $p\Omega$ system, for which a bound state has been reported by the HAL QCD collaboration Etminan et al. (2014); Iritani and others (2019). When the $\Omega$ couples to the $\Xi\bar{K}$ channel, a triangle diagram naturally enters the $p\Omega$ interaction, bearing qualitative similarity to the mechanism discussed for $N\phi$ scattering. A crucial difference is that the $\Xi$ and $\bar{K}$ lines in the triangle diagram are no longer interchangeable. As a result, the ultraviolet divergence in the loop integral cannot be eliminated by symmetry and is instead tied to coupled-channel dynamics Sekihara et al. (2018). This feature leads to a stronger attractive interaction in the $p\Omega$ system than in the $N\phi$ case. In the meantime, the comparison suggests that the attraction generated by a convergent triangle diagram alone is insufficient to form a bound state at an unphysical kaon mass around $500\,\rm{MeV}$.
@@ IV Summary
In summary, a triangle diagram promoted by the near-threshold $\Lambda(1405)$ is introduced to describe the $N\phi$ scattering in the absence of direct spin interactions. The scattering length, estimated from the convergent part of the triangle diagram with a single cutoff parameter $\Lambda$ in the triangle loop integral, is found to be consistent with the lattice QCD results from the HAL QCD collaboration Lyu et al. (2022). Following the analytic expression deduced from the triangle diagram, the real part of the $N\phi$ scattering length evaluated with physical hadron masses shows a clear overlap with the experimental extraction by the ALICE collaboration Acharya and others (2021).
In the low-energy region, the triangle loop integral leads to a power-law dependence of the scattering length on the difference of the $K\bar{K}$ threshold to $\phi$ mass, $\tilde{\delta}^{0,\,1/2}$. This behavior indicates that the interaction driven by the promoted triangle diagram is different from both the Van der Waals force Appelquist and Fischler (1978); Brambilla et al. (2017) and the long-range tail of two-pion exchange Lyu et al. (2022). Such a nontrivial power counting reflects the interplay between two-body and three-body dynamics near threshold.
@@ Acknowledgement
In this appendix, we examine possible contributions from two-pion and kaon exchange to the $N\phi$ interaction. The corresponding Feynman diagrams are shown in Fig. 9. Both $S$- and $P$-wave transitions are in principle allowed. The $S$-wave transition (denoted by the rectangle) is governed by the Weinberg-Tomozawa (WT) term at leading order in chiral perturbation theory, while the $P$-wave transition arises from the axial coupling.
However, the WT interactions for elastic $\phi\pi$, $\phi K$, and $\phi\bar{K}$ scattering vanish at leading order Roca et al. (2005). As a consequence, the contribution from the two-pion exchange diagram shown in Fig. 9(a) identically vanishes. This situation is in contrast to the case of two-pion exchange in $J/\psi N$ scattering, where a nonvanishing $J/\psi\pi\to J/\psi\pi$ transition exists, as illustrated in Fig. 1(c) of Ref. Tarrús Castellà and Krein (2018). Similarly, the kaon-exchange contribution shown in Fig. 9(b) also vanishes at this order. This again differs from the mechanism in Fig. 1(d) of Ref. Tarrús Castellà and Krein (2018), where the corresponding meson-hadron transition is allowed. Therefore, neither two-pion nor kaon exchange generates a long-range interaction in $N\phi$ scattering in two-point loop. These support the conclusion drawn in the main text: the $N\phi$ interaction near threshold is not dominated by conventional one- or two-meson exchange mechanisms, but is instead driven by the promoted triangle diagram involving near-on-shell intermediate states.
@@
Furthermore, there exists a clear separation of scales between the two mechanisms. The effective ranges associated with two-pion and two-kaon exchange are of order $1/(2m_{\pi})$ and $1/(2m_{K})$, respectively. When the kaons in Fig. 9(d) are nearly on shell, the typical momentum $\sqrt{m_{K}(m_{\phi}-2m_{K})}$ introduces a new dynamical scale in $N\phi$ scattering. We focus on interactions with a range of order $1/\sqrt{m_{K}(m_{\phi}-2m_{K})}$, for which the two-kaon exchange mechanism is expected to be particularly relevant compared with the two-pion exchange.

### v2

@@ III Results and discussion
Kinematical consistency requires the parameter $\Lambda$ in the loop integral to be larger than the typical momentum flowing through the Kaon lines. For a small value, $\Lambda=120\,\mathrm{MeV}$, noticeable deviations from the uncertainty bands are observed. In contrast, for $\Lambda=400$-$1000\,\mathrm{MeV}$ the real part of the $N\phi$ scattering length remains stable and overlaps with the bands, supporting the relevance of the triangle mechanism in the physical-mass case.
In Fig. 8, the imaginary part of $N\phi$ scattering length determined by the open channel effect, should be free of the $\Lambda$ in the triangle. However, the open channel effect is not evaluated directly here and approximated by the $\Lambda(1405)$ pole. When $\Lambda$ is much larger than $|b|$, the green lines show little dependence of $\Lambda$ and the dotted line overlaps with the one extracted by ALICE Acharya and others (2021), which indicates that the triangle diagram is non-negligible in $N\phi$ scattering.
@@
Figure 7: The real part of the $N\phi$ scattering length from Two-Kaon-Exchange with physical masses, where the solid, dashed, and dotted lines correspond to the central value, lower and upper limits of $g_{i}$ and the corresponding $\sqrt{s}_{R}$, respectively.
Figure 8: The imaginary part of the $N\phi$ scattering length from Two-Kaon-Exchange with physical masses, where the solid, dashed, and dotted lines correspond to the central value, lower and upper limits of $g_{i}$ and the corresponding $\sqrt{s}_{R}$, respectively.
@@
Such threshold-enhanced triangle dynamics is not unique to the $N\phi$ system and has been widely discussed in heavy exotic hadrons. Representative examples include the role of $D\bar{D}$ rescattering in $D\bar{D}\pi$ loops for the $X(3872)$ Dai et al. (2020); Yan et al. (2022a); Cao et al. (2024), the contribution of $\Sigma_{c}\bar{D}\pi$ loops to positive-parity $P_{c}(4457)$ states Peng et al. (2021a); Wu et al. (2024b), and the importance of $DD\pi$ loops in the description of the $T_{cc}$ state Fleming et al. (2021); Yan and Valderrama (2022); Jia et al. (2023a, b); Dai et al. (2023). In these systems, the proximity of two- and three-body thresholds leads to non-perturbative enhancements analogous to those identified here.
A similar mechanism is expected to operate in channels where $D\bar{D}\pi$ couples to $Z_{c}(4020)$ or $X(4020)$, promoted by the nearby $X(3872)$ or $W_{c1}(3872)$ poles Peng et al. (2023a); Ji et al. (2025). In such cases, the triangle diagram generates an interaction that plays a role comparable to that of scalar-meson exchange, supplementing the axial-vector exchange mechanism Yan et al. (2021). Related threshold effects have also been discussed for hidden-charm and hidden-strangeness pentaquark states, where the $P_{c}(4457)$ and $P_{cs}(4560)$ states are enhanced by lower-lying $P_{c}(4312)$ and $P_{cs}(4338)$ poles through $\bar{D}\Sigma_{c}\pi$ and $\bar{D}\Xi_{c}\pi$ loops, respectively Peng et al. (2023b); Du et al. (2020); Fernández-Ramírez et al. (2019); Wang and Wang (2020); Liu et al. (2019); Wang et al. (2020); Chen et al. (2021); Yan et al. (2022b); Xiao et al. (2021); Yan et al. (2023); Nakamura and Wu (2023); Zhu et al. (2023); Yang et al. (2022).
@@
In the dibaryon sector, a closely related dynamical mechanism has been invoked to interpret the $d(2380)$ as a $\Delta\Delta$ molecular state, where the existence of a near-threshold $N\Delta$ bound state plays a crucial role Gal and Garcilazo (2013). A similar picture also emerges in the $IJ=21$ and 1212 $N\Delta$ scattering channels when poles in the $NN$ subsystem are taken into account Gal and Garcilazo (2014). However, the situation for the $d(2380)$ is considerably more involved: the resonance is located between the $N\Delta\pi$ and $\Delta\Delta$ thresholds, and the large intrinsic width of the $\Delta$ renders a clean separation between two-body and three-body dynamics highly nontrivial Peng et al. (2022). This contrasts with the $N\phi$ system, where the thresholds are much better separated, and the triangle mechanism can be isolated more transparently.
Another instructive comparison is provided by the $p\Omega$ system, for which a bound state has been reported by the HAL QCD collaboration Etminan et al. (2014); Iritani and others (2019). When the $\Omega$ couples to the $\Xi\bar{K}$ channel, a triangle diagram naturally enters the $p\Omega$ interaction, bearing qualitative similarity to the mechanism discussed for $N\phi$ scattering. A crucial difference is that $m_{\Xi}+m_{\bar{K}}-m_{\Omega}=143\,\rm{MeV}$ leads to a relativistic Kaon in the triangle diagram and a different loop integral over $l^{0}$ in Eq. (8). As a result, the interaction differs from the one in Eq. (15) and is instead tied to coupled-channel dynamics Sekihara et al. (2018). This feature leads to a stronger attractive interaction in the $p\Omega$ system than in the $N\phi$ case. In the meantime, the comparison suggests that the attraction generated by a convergent triangle diagram alone is insufficient to form a bound state at an unphysical Kaon mass around $500\,\rm{MeV}$.
@@ IV Summary
In summary, we investigate low-energy $N\phi$ scattering via a triangle diagram promoted by the near-threshold $\Lambda(1405)$ pole, in the absence of direct spin interactions. The scattering length, evaluated from the triangle loop integral, is found to be consistent with the lattice QCD results from the HAL QCD collaboration Lyu et al. (2022). Furthermore, using the analytic expression derived from the triangle diagram with physical hadron masses, the obtained $N\phi$ scattering length shows a clear overlap with the experimental extraction by the ALICE collaboration Acharya and others (2021).
In the low-energy region, the triangle loop integral relies on the $\Lambda(1405)$. This behavior indicates that the interaction driven by the promoted triangle diagram is different from both the Van der Waals force Appelquist and Fischler (1978); Brambilla et al. (2017) and the long-range tail of two-pion exchange Lyu et al. (2022).
@@ Acknowledgement
In this appendix, we examine possible contributions from two-pion and Kaon exchange to the $N\phi$ interaction. The corresponding Feynman diagrams are shown in Fig. 9. Both $S$- and $P$-wave transitions are in principle allowed. The $S$-wave transition is governed by the Weinberg-Tomozawa (WT) term at leading order in chiral perturbation theory, while the $P$-wave transition arises from the axial coupling.
However, the WT interactions for elastic $\phi\pi$, $\phi K$, and $\phi\bar{K}$ scattering vanish at leading order Roca et al. (2005). As a consequence, the contribution from the two-pion exchange diagram shown in Fig. 9(a) identically vanishes. This situation is in contrast to the case of two-pion exchange in $J/\psi N$ scattering, where a nonvanishing $J/\psi\pi\to J/\psi\pi$ transition exists, as illustrated in Fig. 1(c) of Ref. Tarrús Castellà and Krein (2018). Similarly, the Kaon-exchange contribution shown in Fig. 9(b) also vanishes at this order. This again differs from the mechanism in Fig. 1(d) of Ref. Tarrús Castellà and Krein (2018), where the corresponding meson-hadron transition is allowed. Therefore, neither two-pion nor Kaon exchange generates a long-range interaction in $N\phi$ scattering in two-point loop. These support the conclusion drawn in the main text: the $N\phi$ interaction near threshold is not primarily a result of conventional one- or two-meson exchange mechanisms, but is instead driven by the promoted triangle diagram involving near-on-shell intermediate states.
@@
Furthermore, the proposed mechanism is distinct from the conventional two-pion exchange in kinematics, dynamics, and its absorptive nature. Kinematically, there is a clear separation of scales. The effective range of the two-pion exchange is governed by the hard scale $1/(2m_{\pi})$. In contrast, the near-on-shell propagation of the intermediate Kaons introduces new soft dynamical scales, characterized by the momenta $|a_{0}|=\sqrt{-m_{K}\delta}$ and $|b|=\sqrt{-2m_{\Lambda}\Delta E}$. We focus on interactions within the range of order $1/|b|$, where the long-range physics of the triangle diagram dominates. Dynamically, the two mechanisms differ fundamentally. The two-pion exchange is suppressed in the low-energy region due to the vanishing leading-order Weinberg-Tomozawa interaction in the $\phi\pi\to\phi\pi$ channel. Conversely, the two-Kaon exchange is significantly promoted by the non-perturbative three-body dynamics involving the near-threshold $\Lambda(1405)$ pole. Crucially, unlike the real-valued potential often assumed in simple exchange models, the two-Kaon mechanism naturally generates an imaginary part for the $N\phi$ scattering length due to the open channel effects and the complex pole structure. As presented in the Fig. 8, our calculated imaginary part is in good agreement with the femtoscopic extraction by the ALICE collaboration. This consistency with experimental data provides strong support for the dominance of the pole-enhanced triangle mechanism over the suppressed two-pion exchange in the near-threshold region.